# Belobog: Move Language Fuzzing Framework For Real-World Smart Contracts

### v1

@@ Abstract.
Move is a research-oriented programming language design for secure and verifiable smart contract development and has been widely used in managing billions of digital assets in blockchains, such as Sui and Aptos. Move features a strong static type system and explicit resource semantics to enforce safety properties such as the prevention of data races, invalid asset transfers, and entry vulnerabilities. However, smart contracts written in Move may still contain certain vulnerabilities that are beyond the reach of its type system. It is thus essential to validate Move smart contracts. Unfortunately, due to its strong type system, existing smart contract fuzzers are ineffective in producing syntactically or semantically valid transactions to test Move smart contracts.
@@
We evaluated Belobog on 109 real-world Move smart contract projects. The experimental results show that Belobog is able to detect 100% critical and 79% major vulnerabilities manually audited by human experts. We further selected two recent notorious incidents in Move smart contracts, i.e., Cetus and Nemo. Belobog successfully reproduced full exploits for both of them, without any prior knowledge.
@@ 1. Introduction
Smart contracts have become a core component of modern blockchain systems, enabling decentralized applications to automate interactions. Since smart contracts have been used in managing numerous digital assets, their correctness and security are critical. Bugs in deployed smart contracts can result in severe consequences, such as financial loss (zhou2023sok). The most prevailing smart contract languages, Solidity (solidity), adopt an easy-to-use language design but expose developers to subtle security pitfalls, such as reentrancy attacks. These recurring issues highlight the need for languages that provide stronger security guarantees by design.
The Move language emerges as a new smart contract language design with strong type safety. Its static type system and resource-oriented semantics enforce key safety properties at both compile and run time, preventing common bugs like reentrancy. More and more mainstream blockchains have chosen Move as their language for developing smart contracts, such as Sui(suimove) and Aptos(aptosmove). Despite its strong type system, smart contracts written in Move can still contain logic flaws or other bugs that are beyond the scope of the type system, such as the notorious bugs in Cetus (cetusreport) and Nemo (nemoreport). Therefore, it is still essential to test and verify a Move smart contract. Unfortunately, existing smart contract fuzzers are largely designed for the Ethereum Virtual Machine (EVM)(evm) and lack awareness of Move’s unique type system. As our evaluation in Section 2.1 will show, they all fail to generate enough valid transactions — most generated transactions are rejected by the Move Virtual Machine before reaching meaningful code paths.
@@
Challenge 1: Type safety of objects. Move represents on-chain states as strongly-typed objects, which results in several strict rules for objects, i.e., (1) objects that are passed as function inputs must match the function’s signatures, (2) objects can only be created in the module where they were defined, and (3) a function caller can only use its owned or shared objects to interact with other functions. Thus, fuzzers cannot fill random bytes to forge objects but must instead construct them through legitimate function calls or reuse existing ones. For example, the second argument type of repay<T>() in Figure 1(a) is Receipt<T>, which is defined in the same module as repay<T>(). In this module, the only function that can produce objects of type Receipt<T> is loan<T>(), and thus a valid transaction must call loan<T>() first to create a usable object of type Receipt<T> before calling loan<T>().
@@
Challenge 3: Type abilities of objects. All Move objects can have several associated abilities to constrain how they can be used and persisted. For example, the struct type VeryAble in Figure 1(a) has drop and store abilities while Receipt has none of them. The drop and store are two abilities that may affect the validity of a transaction. The store ability determines if an object can be stored in persistent storage, while the drop ability allows an object to be dropped without being used. An object without drop and store abilities, such as Receipt, can not be transferred to other accounts and thus must be consumed by one and only one function call within the same transaction. Such objects are usually called “Hot Potato” and are typically used for implementing flash loans (flashloan). It enables users to perform permissionless and collateral-free loans and ensures that the loans are repaid within a single transaction. When generating a transaction, we have to guarantee that all produced “Hot Potato” objects are properly handled. For example, if we create a function call loan<T>() in a transaction, then the created object of type Receipt<T> should be passed as input to another function to consume it. Such “Hot Potato” objects introduce constraints across different function calls.
Our contributions. In this paper, we propose a type graph-guided Move fuzzing framework, Belobog. Given a Move smart contract, we build a type graph to model all the type constraints. This type graph will later be used to guide the generation and mutation of transactions during fuzzing. For example, Figure 1(b) shows a simplified type graph for the smart contract in Figure 1(a). Nodes in the graph represent different object types or functions, while the annotated edges maintain different relations and constraints among them. In order to further overcome tricky checks in Move smart contracts, we also design and implement a concolic executor in Belobog.
We extensively evaluate Belobog on a large-scale dataset provided by a well-known Web3 audit company, Movebit. and a real-world dataset collected by an empirical work (movescan). The results are very promising: Belobog is able to detect 100% critical vulnerabilities and 79% major vulnerabilities when cross-compared with manually audited results by human experts. We also selected two notorious incidents in the Move ecosystem, Cetus with \$200 million loss and Nemo with \$3 million loss, and Belobog successfully reproduced both incidents with full exploits. We envision that this work bridges the gap between Move’s static safety guarantees and dynamic testing and provides a general framework for improving the overall security and reliability of Move‑like smart contract ecosystems. We have open-sourced Belobog at https://github.com/abortfuzz/belobog.
@@ 2. Motivation
In this section, we motivate our design by first showing the deficiency of existing fuzzers on Move smart contracts, then empirically demonstrating the criticality of handling the special constraints in Move, and lastly examplifying how Belobog can generate valid transactions with the guidance of type graph.
@@ 2.1. Deficiency of Existing Fuzzing Tools on Move
ItyFuzz and Sui-Fuzzer are two state-of-the-art smart contract fuzzers that can be used on Move smart contracts. We create a Move smart contract with the example shown in Figure 1(a). This example, although constructed, includes the main and commonly seen features in Move smart contracts. We try to run both ItyFuzz and Sui-Fuzzer on this smart contract. Unfortunately, all the transactions generated by them have type-mismatch errors and are rejected without execution, such as the one shown in Figure 2(a). The root cause is that neither of them supports type parameters, which are used in the structs and functions in our example.
We further simplify the smart contract by removing all the type parameters and the type Coin<T> from a standard library. However, Sui-Fuzzer still fails to generate any valid transactions, as shown in Figure 2(b). The reason is two-fold. First, it will internally forge objects by filling random bytes, which can easily violate the type safety rule of Move. Second, it is not aware of the type abilities of objects and thus “Hot Potato” objects are not properly handled. ItyFuzz generates some valid transactions in this simplified smart contract by randomly mutating function calls until all “Hot Potato” objects are handled. This random mechanism is highly inefficient and can only produce “Hot Potato” objects without type parameters. Although existing fuzzers fail to generate valid transactions on smart contracts with type parameters and “Hot Potato”, we still lack understanding of how common such features are in the real world. In order to measure the prevalence of type parameters and “Hot Potato”, we collected all two billion transactions on the Sui blockchain from November 2024 to November 2025. Since Move smart contracts are organized as packages, we obtain the source packages that each transaction uses. For all the packages, we count the number of packages having type parameters and “Hot Potato”. Figure 3 shows the percentage of packages that have type parameters, “Hot Potato”, and at least one of them. Packages are ranked based on the number of Move function calls that use them. The result shows that among the most used (90%) packages, more than 80% of them have at least type parameter or “Hot Potato”, indicating the prevalence of such Move features in real-world.
@@
(c) Belobog successfully generate a valid transaction by correctly instantiating type parameters and handling “Hot Potato”.
@@ 2.2. Effectiveness of Belobog
Belobog overcomes the limitations of previous fuzzers by implementing type-aware transaction generation and mutation. With the statically built type graph for a smart contract, Belobog can always generate valid transactions. For example, to generate a transaction for Figure 1(a), Belobog first queries the type graph to understand that the output object of loan<T>() contains a “Hot Potato”, which can be luckly consumed by repay<K>(). Since there are no additional type constraints, the two type parameters T and K can be instantiated with the same concrete type USDC. With all the information obtained from the type graph, Belobog can always produce valid transactions for testing this smart contract, as shown in Figure 2(c).
@@ 3.1. Preliminaries
An object refers to a typed entity. Each object is an instance of a Move struct or enum type that possesses the key abilities.
Objects can be created by a transaction or persist in on-chain blockchain storage. The v1 and v2 in Figure 2(c) are two objects created by the transaction. On-chain objects are created and stored by some transactions. A new transaction can use on-chain objects owned by the transaction sender, or are simply shared objects. On-chain objects are non-“Hot Potato” by definition and thus a fuzzer can directly use them as long as the type is matched.
@@
A transaction $S$ in Move is a sequence of function calls without control flows.
On both Sui and Aptos, the two prevailing blockchains that adopt the Move language, a transaction could carry multiple Move function calls with slightly different implementation details. Sui supports Programmable Transaction Block (PTBs) composed with a sequence of function calls, each of which can use either an existing object or a value returned from previous Move function calls. Aptos allows users to ad-hoc compile and run a piece of Move code as a transaction, named Script, instead of deploying a Move contract, which essentially is a superset of Programmable Transaction Block with arbitrary control flows. Both mechanisms enable a transaction to do multiple Move function calls in an atomic fashion. Thus, without loss of generality, transaction in this paper is defined as a sequence of function calls without control flows.
@@ 3.2. Belobog in a Nutshell
  • Step 2: Type graph construction. Since Move is a strongly typed language, we need to ensure the type validity of all the generated/mutated transactions. To this end, we construct a type graph $G$ to capture all necessary type-related information, which can then be used to guide the generation and mutation of transactions. Our usage of the type graph $G$ ensures that all the transactions produced are well-typed. Technical details will be introduced in Section 3.3.
@@
      For random generation, Belobog first selects a random function node in the type graph. Then, Belobog walks on the type graph to find additional function nodes that are required to make a valid transaction. For each function call, Belobog guarantees that each input object is produced by some earlier function or from $\texttt{Pool}_{obj}$, and each output “Hot Potato” object is properly consumed.
@@
      For random mutation, Belobog first selects a seed transaction from the seed pool and then mutates it by modifying constant values or adding/removing function calls. Since newly added or removed function calls may break type validity, Belobog again uses the type graph to maintain validity.
@@
      Concolic execution. The above generation and mutation processes only care about the type validity of a transaction. For primitive types such as u64 that are used as function inputs, they all use a random strategy to choose or mutate their values. In fact, hard-coded conditions or checks are prevalent in smart contracts. For example, line LABEL:line:fee in Figure 4(a) can only be triggered when the inputs satisfy the value constraint, which is nearly impossible to reach with random values. To tackle this challenge, we design and implement a concolic executor based on the Move Virtual Machine.
    Technical details about transaction generation and mutation will be introduced in Section 3.4. Technical details about the concolic executor will be introduced in Section 3.5.
  • Step 4: Transaction execution and feedback collection. The generated new transaction will then be executed on the Move Virtual Machine (MoveVM). If the transaction increases the code coverage [^1] We use the standard branch coverage in our implementation., it will be added to the seed corpus for further mutation. If the execution reveals an oracle violation, it will be reported as a potential vulnerability. Different from C/C++ software, which typically has a standard oracle, such as buffer overflow, there is no standard oracle in smart contracts, and the oracle problem remains an unsolved problem (defiranger; verite; ityfuzz). In our implementation, we provide five pre-defined common oracles that we observe in Move smart contracts. Additionally, we expose a user-defined oracle interface that enables the tool users to define their own oracles. Technical details about oracle definition will be discussed in Section 4.
@@ 3.3. Type Graph Construction
Move smart contracts usually rely on standard libraries, or third-party smart contracts. Consequently, a smart contract may use types defined somewhere else. In order to build a type graph with complete information, we need to traverse all the libraries and smart contracts that are referenced by our target smart contract. In our implementation, Belobog supports building the type graph for a group of smart contracts at the same time.
Our type graph is inspired by the signature graph introduced in program synthesis (signaturegraph) and the API graph used in API-driven program synthesis (apigraph). Our type graph construction initializes the type graph $G$ with every used primitive type as a default type node $n_{t}$. Then, it analyzes all the struct types. If the struct has at least drop or store ability, it is added as a default type node $n_{t}$, illustrated as ovals in Figure 4(b); Otherwise, it is added as a “Hot Potato” type node $n_{h}$, illustrated as double-circled ovals in Figure 4(b). Then, for each function definition $f$ in the target smart contract, we proceed as follows.
@@
Figure 5. Illustrative transaction generation starting from a graph trace (left), then concretizing type parameters by type substitution (middle), and finally instantiating a transaction (right). The dashed red nodes demonstrate the mutation process of graph trace extension.

### v2

@@ Abstract.
Move is a research-oriented programming language designed for secure and verifiable smart contract development and has been widely used in managing billions of digital assets in blockchains, such as Sui and Aptos. Move features a strong static type system and explicit resource semantics to enforce safety properties such as the prevention of data races, invalid asset transfers, and entry vulnerabilities. However, smart contracts written in Move may still contain certain vulnerabilities that are beyond the reach of its type system. It is thus essential to validate Move smart contracts. Unfortunately, due to its strong type system, existing smart contract fuzzers are ineffective in producing syntactically or semantically valid transactions to test Move smart contracts.
@@
We evaluated Belobog on 109 real-world Move smart contract projects. The experimental results show that Belobog is able to detect 100% critical and 79% major vulnerabilities manually audited by human experts. We further selected two recent notorious incidents in the Move ecosystem, i.e., Cetus and Nemo. Belobog successfully reproduced full exploits for both of them, without any prior knowledge. Moreover, we applied Belobog on three ongoing auditing projects and found 2 critical, 2 major, and 3 medium new vulnerabilities, all acknowledged by the project developers.
@@ 1. Introduction
Smart contracts have become a core component of modern blockchain systems, enabling decentralized applications to automate interactions. Since smart contracts have been used to manage numerous digital assets, their correctness and security are critical. Bugs in deployed smart contracts can result in severe consequences, such as financial loss (Zhou et al., 2023). The most prevailing smart contract language, Solidity (24), adopt an easy-to-use language design but expose developers to subtle security pitfalls, such as reentrancy attacks. These recurring issues highlight the need for languages that provide stronger security guarantees by design.
The Move language emerges as a new smart contract language designed with strong type safety. Its static type system and resource-oriented semantics enforce key safety properties at both compile and run time, preventing common bugs like reentrancy. The Move language has been increasingly adopted by mainstream blockchains. For example, Sui(Welc and Blackshear, 2023) and Aptos(1), the two prevailing blockchains with 6.4 billion market capitalization and 1.5 million daily active users (2), use the Move language for development. Despite its strong type system, smart contracts written in Move can still contain logic flaws or other bugs that are beyond the scope of the type system, such as the notorious bugs in Cetus (3) and Nemo (18), causing millions of dollars of loss. Therefore, it is still essential to test and verify a Move smart contract. Unfortunately, existing smart contract fuzzers are largely designed for the Ethereum Virtual Machine (EVM)(Wood, 2024) and lack awareness of Move’s unique type system. As our evaluation in Section 2.1 will show, they all fail to generate enough valid transactions—most transactions are rejected by the Move Virtual Machine before reaching meaningful code paths.
@@
Challenge 1: Type safety of objects. Move represents on-chain states as strongly-typed objects, which results in several strict rules for objects, i.e., (1) objects that are passed as function inputs must match the function’s signatures, (2) objects can only be created in the module where they were defined, and (3) a function caller can only use its owned or public shared objects to interact with other functions. Thus, fuzzers cannot fill random bytes to forge objects but must instead construct them through legitimate function calls or reuse existing ones. For example, the second argument type of repay<T>() in Figure 1(a) is Receipt<T>, which is defined in the same module as repay<T>(). In this module, the only function that can produce objects of type Receipt<T> is loan<T>(), and thus a valid transaction must call loan<T>() first to create a usable object of type Receipt<T> before calling repay<T>().
@@
Challenge 3: Type abilities of objects. All Move objects can have several associated abilities to constrain how they can be used and persisted. For example, the struct type VeryAble in Figure 1(a) has drop and store abilities while Receipt has none of them. The drop and store are two abilities that may affect the validity of a transaction. The store ability determines if an object can be stored in persistent storage, while the drop ability allows an object to be dropped without being used. An object without drop and store abilities, such as Receipt, can not be transferred to other accounts and thus must be consumed by one and only one function call within the same transaction. Such objects are usually called “Hot Potato” and are typically used for implementing flash loans (Qin et al., 2021). It enables users to perform permissionless and collateral-free loans and ensures that the loans are repaid in a single transaction, where a transaction on Move refers to a sequence of Move function calls as our Section 3.1 will explain in detail. When generating a transaction, a fuzzer must ensure that all produced “Hot Potato” objects are properly handled. For example, if a fuzzer creates a function call loan<T>() in a transaction, then the created object of type Receipt<T> should be passed as input to another function, in this case, only repay<T> to consume it. Such “Hot Potato” objects introduce constraints across different function calls.
Our contributions. In this paper, we propose a type graph-guided Move fuzzing framework, Belobog. Given a Move smart contract, we build a type graph to model all the type constraints. The type graph will later be used to guide the generation and mutation of transactions during fuzzing. For example, Figure 1(b) shows a simplified type graph for the smart contract in Figure 1(a). Nodes in the graph represent different object types or functions, while the annotated edges maintain different relations and constraints among them. In order to further overcome tricky checks in Move smart contracts, we also design and implement a concolic executor in Belobog.
We extensively evaluate Belobog on a large-scale dataset provided by a well-known Web3 audit company and a real-world dataset collected by an empirical work (Song et al., 2024). The results are very promising: Belobog is able to detect 100% critical vulnerabilities and 79% major vulnerabilities when cross-compared with manually audited results by human experts. We also selected two notorious incidents in the Move ecosystem, Cetus with \$200 million loss and Nemo with \$2.6 million loss. Belobog successfully reproduced both incidents with full exploits. We envision that this work bridges the gap between Move’s static safety guarantees and dynamic testing.
@@ 2. Motivation
In this section, we motivate our design by first showing the deficiency of existing fuzzers on Move smart contracts, then empirically demonstrating the criticality of handling the special constraints in Move, and lastly exemplifying how Belobog can generate valid transactions with the guidance of the type graph.
@@ 2.1. Deficiency of Existing Fuzzing Tools on Move
ItyFuzz and Sui-Fuzzer are two state-of-the-art smart contract fuzzers that can be used on Move smart contracts. To illustrate their deficiency, we create a Move smart contract with the example shown in Figure 1(a). This example, although constructed, includes the main and commonly seen features in Move smart contracts. We try to run both ItyFuzz and Sui-Fuzzer on this smart contract. Unfortunately, all the transactions generated by them have type-mismatch errors and thus are rejected without execution, such as the one shown in Figure 2(a). The root cause is that neither of them supports type parameters, which are used in the structs and functions in our example.
We further simplify the smart contract by removing all the type parameters and the type Coin<T> from the standard library. However, Sui-Fuzzer still fails to generate any valid transactions, as shown in Figure 2(b). The reason is two-fold. First, it will internally forge objects by filling random bytes, which can easily violate the type safety rule of Move. Second, it is not aware of the type abilities of objects and thus “Hot Potato” objects are not properly handled. ItyFuzz generates some valid transactions in this simplified smart contract by randomly mutating function calls until all “Hot Potato” objects are handled. This random mechanism is highly inefficient and can only produce “Hot Potato” objects without type parameters. Although existing fuzzers fail to generate valid transactions on smart contracts with type parameters and “Hot Potato”, we still lack an understanding of how common such features are in the real world. In order to measure the prevalence of type parameters and “Hot Potato”, we collected all two billion transactions on the Sui blockchain from November 2024 to November 2025. Since Move smart contracts are organized as packages, we obtain the source packages that each transaction uses. For all the packages, we count the number of packages with type parameters and “Hot Potato”. Figure 3 shows the percentage of packages that have type parameters, “Hot Potato”, and at least one of them. Packages are ranked based on the number of Move function calls that use them. The result shows that among the top 90% most-used packages, more than 80% of them have at least one type parameter or “Hot Potato”, indicating the prevalence of such Move features in the real-world deployment.
@@
(c) Belobog successfully generates a valid transaction by correctly instantiating type parameters and handling “Hot Potato”.
@@ 2.2. Effectiveness of Belobog
Belobog overcomes the limitations of previous fuzzers by implementing type-aware transaction generation and mutation. With the statically built type graph for a smart contract, Belobog can always generate valid transactions. For example, to generate a transaction for Figure 1(a), Belobog first queries the type graph to understand that the output object of loan<T>() contains a “Hot Potato” that can be consumed by repay<K>(). Since there are no additional type constraints, the two type parameters T and K can be instantiated with the same concrete type USDC. With all the information obtained from the type graph, Belobog can always produce valid transactions for testing this smart contract, as shown in Figure 2(c).
@@ 3.1. Preliminaries
An object refers to a typed entity. Each object is an instance of a Move struct or enum type that possesses certain abilities.
Objects can be created by a transaction or persisted in on-chain blockchain storage. The v1 and v2 in Figure 2(c) are two objects created by the transaction. On-chain objects are created and stored by some transactions. A new transaction can use on-chain objects owned by the transaction sender or shared objects. On-chain objects are non-“Hot Potato” by definition, and thus a fuzzer can directly use them as long as the type is matched.
@@
A transaction $S$ in Move is a sequence of function calls without control flow, where each function call may use primitive values or existing objects as parameters.
On both Sui and Aptos, the two prevailing blockchains that adopt the Move language, a transaction can carry multiple Move function calls with slightly different implementation details. Sui supports Programmable Transaction Block (PTBs) composed of a sequence of function calls, each of which can use either an existing object or a value returned from previous Move function calls. Aptos allows users to ad-hoc compile and run a piece of Move code as a transaction, named Script, instead of deploying a Move contract, which is essentially a superset of Programmable Transaction Block with arbitrary control flows. Both mechanisms enable a transaction to do multiple Move function calls in an atomic fashion. Thus, without loss of generality, a transaction in this paper is defined as a sequence of function calls without control flows.
@@ 3.2. Belobog in a Nutshell
  • Step 2: Type graph construction. Since Move is a strongly typed language, we need to ensure the type validity of all the generated/mutated transactions. To this end, we construct a type graph $G$ to capture all necessary type-related information, which can then be used to guide the generation and mutation of transactions. Our usage of the type graph $G$ ensures that all the transactions produced are well-typed. (Details in Section 3.3)
@@
      For random generation, Belobog first selects a random function node in the type graph. Then, Belobog walks on the type graph to find additional function nodes that are required to make a valid transaction. For each function call, Belobog guarantees that each input object is produced by some earlier function or from $\texttt{Pool}_{obj}$, and each output “Hot Potato” object is properly consumed. (Details in Section 3.4)
@@
      For random mutation, Belobog first selects a seed transaction from the seed pool and then mutates it by modifying constant values or adding/removing function calls. Since newly added or removed function calls may break type validity, Belobog again uses the type graph to maintain validity. (Details in Section 3.4)
@@
      Concolic execution. The above generation and mutation processes only care about the type validity of a transaction. For primitive types such as u64 that are used as function inputs, they all choose or mutate their values randomly. However, such a random strategy is hard to cover hard-coded conditions or checks that are prevalent in smart contracts. For example, line LABEL:line:fee in Figure 4(a) can only be triggered when the inputs satisfy the value constraint, which is nearly impossible to reach with random values. To tackle this challenge, we design and implement a concolic executor based on the Move Virtual Machine. (Details in Section 3.5 and our supplementary materials)
  • Step 4: Transaction execution and feedback collection. The generated new transaction will then be executed on the Move Virtual Machine (MoveVM). If the transaction increases the code coverage [^1] We use the standard branch coverage in our implementation., it will be added to the seed corpus for further mutation. If the execution reveals an oracle violation, it will be reported as a potential vulnerability. Different from C/C++ software, which typically has a standard oracle, such as buffer overflow, there is no standard oracle in smart contracts, and the oracle problem remains an unsolved problem (Wu et al., 2023; Kong et al., 2025; Shou et al., 2023). In our implementation, we provide five pre-defined common oracles that we observe in Move smart contracts. Additionally, we expose a user-defined oracle interface that enables the tool users to define their own oracles. (Details in Section 4)
@@ 3.3. Type Graph Construction
Move smart contracts usually rely on standard libraries or third-party smart contracts. Consequently, a smart contract may use types defined somewhere else. In order to build a type graph with complete information, we need to traverse all the libraries and smart contracts that are referenced by our target smart contract. In our implementation, Belobog supports building the type graph for a group of smart contracts at the same time.
Our type graph is inspired by the signature graph introduced in program synthesis (Mandelin et al., 2005) and the API graph used in API-driven program synthesis (Sotiropoulos et al., 2024). Our type graph construction initializes the type graph $G$ with every used primitive type as a default type node $n_{t}$. Then, it analyzes all the struct types. If the struct has at least drop or store ability, it is added as a default type node $n_{t}$, illustrated as ovals in Figure 4(b); Otherwise, it is added as a “Hot Potato” type node $n_{h}$, illustrated as double-circled ovals in Figure 4(b). Then, for each function definition $f$ in the target smart contract, we proceed as follows.
@@
Figure 5. Illustrative transaction generation starting from a graph trace (left), then concretizing type parameters by type substitution (middle), and finally instantiating a transaction (right). The dashed nodes demonstrate the mutation process of graph trace extension.